OpenStudy (anonymous):

1.6x-3=-2(4.2x-6) solve for x

OpenStudy (anonymous):

\[1.6x-3=-2(4.2x-6)\]\[1.6x-3=-8.4x-12\]\[6.8x=-9\]\[x=45/34=1.3235\]

OpenStudy (anonymous):

thank you!!!!

OpenStudy (anonymous):

No problem :)

razor99 (razor99):

thats worng!!!!!!!!!

Parth (parthkohli):

It'll be -8.4x + 12

OpenStudy (anonymous):

it is then what is the right answer

razor99 (razor99):

it should be x=3/2

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow 1.6x - 3 = -8.4x + 12 }\) \(\Large \color{Black}{\Rightarrow 1.6x = -8.4x + 15 }\) \(\Large \color{Black}{\Rightarrow 10x = 15 }\) \(\Large \color{Black}{\Rightarrow x = {15 \over 10} = 1.5 }\) Here's the correct answer.

Parth (parthkohli):

@megan1124 Got it?

OpenStudy (anonymous):

oops yeah. parthkohli's is correct. sorry

OpenStudy (anonymous):

No wait!!

OpenStudy (anonymous):

why is it 10x?

OpenStudy (anonymous):

its ok thats what i put anyway because its multiple choice and that was the closest one to your answer :)

OpenStudy (anonymous):

Oh, yeah got it haha

OpenStudy (anonymous):

forget what i said.

razor99 (razor99):

@ParthKohli your answer is wrong tooo!!!!!!!!!!!

OpenStudy (anonymous):

why is it wrong ur wrong razor

Parth (parthkohli):

3/2 = 1.5 FAIL lol

Parth (parthkohli):

3/2 is same as 1.5 @megan1124

OpenStudy (anonymous):

ok that is what i thought

OpenStudy (anonymous):

I agree with razor99 substitute the answer instead of x and check whether they are the same. Thats my opinion

razor99 (razor99):

see i told u guyz

OpenStudy (anonymous):

it is because its multiple choice and the answers are negative

OpenStudy (anonymous):

either way x= 1.323529412

OpenStudy (anonymous):

yes but that is not an answer

OpenStudy (anonymous):

what are the options?

Parth (parthkohli):

Then just plug in.

OpenStudy (anonymous):

a.1.5 b. is wrong because my teacher said so c. -.5 d. -.9

OpenStudy (anonymous):

what's b?

OpenStudy (anonymous):

1.2 but its wrong

OpenStudy (anonymous):

It has to be 1.5 http://www.wolframalpha.com/input/?i=1.6x-3%3D-2%284.2x-6%29

OpenStudy (anonymous):

thanks :)